# Data Readiness for AI: A 360-Degree Survey

### v1

@@ 4.2.1.5 Bias Indicator
Example: Consider a dataset of job application essays that will be used to train an AI system to evaluate and rank applicants. A bias indicator can be employed to analyze the textual data, ensuring fairness and reducing bias. The indicator would examine language and semantic patterns to identify potential biases, such as gender, race, or age-related biases. For instance, if the indicator detects a bias where certain occupations or characteristics are consistently favored or discriminated against, it would flag it for further examination. This enables developers to address and mitigate biases in the dataset before training the AI model, promoting equal opportunities and minimizing discriminatory outcomes during the evaluation process. The bias indicator plays a crucial role in preparing the textual data, allowing the creation of more objective and unbiased AI systems for job application assessments.
Metrics in Literature: Papakyriakopoulos et al. (Papakyriakopoulos et al., 2020) propose a robust bias measurement technique using word embeddings and cosine similarity calculations. The method involves defining word pairs to represent different types of discrimination and creating a list of concepts for measuring bias. For variable concepts, which change based on social groups, the bias calculation equation considers the cosine similarity between word pair embeddings and concept embeddings. A modified bias calculation equation is used for non-variable concepts, which remain the same irrespective of social groups. By quantifying the differences in cosine distances, the proposed method comprehensively analyzes bias in different contexts, accounting for variable and non-variable concepts.
Summary: A bias indicator is used to quantify and identify potential biases in textual data for AI applications. Papakyriakopoulos et al. introduced a bias indicator that uses word embeddings and cosine similarity calculations.
@@ 4.2.2. Multimedia Data
This section will focus on exploring the key aspects of data readiness for image, speech, and video data. Specifically, we will concentrate on quality evaluation metrics for each data type. The reason for this exclusive emphasis on quality evaluation metrics is that they are fundamentally crucial in determining the reliability and usability of data for AI applications.
Quality evaluation metrics provide a standardized way to quantify multimedia data accuracy, consistency, and overall reliability. By examining and understanding these metrics, we can ascertain the suitability of data for training AI models effectively and achieving reliable results. Moreover, quality metrics play a pivotal role in identifying potential biases, noise, or inconsistencies in the data, which are essential steps in preparing high-quality datasets for AI algorithms.
While other aspects of data readiness, such as preprocessing techniques or data augmentation methods, are valuable in improving the dataset’s performance, quality evaluation metrics are the foundation for data quality assessment. By focusing on quality evaluation metrics, we can comprehensively analyze the readiness of data, thus ensuring a more in-depth exploration and understanding of the integrity of data for AI applications.
@@ 4.2.2.1 Image Quality
As a metric to measure the readiness of image data for AI, image quality refers to the degree to which an image accurately represents the visual content it intends to depict. It includes various aspects such as sharpness, clarity, color accuracy, and the absence of artifacts or distortions. Assessing image quality is crucial in AI applications as it directly impacts the reliability and effectiveness of algorithms that rely on visual data. High-quality images provide a solid foundation for AI to extract meaningful features, recognize patterns, and make accurate decisions.
Example: In an AI system designed for autonomous driving, image quality plays a vital role in ensuring the accuracy and reliability of object detection and recognition. High-quality images with clear details and accurate colors allow the AI system to accurately identify pedestrians, vehicles, and traffic signs, enabling it to make precise decisions in real-time. Conversely, low-quality images with blurriness or artifacts may lead to misinterpretation of objects or false detections, compromising the safety and efficiency of the autonomous driving system.
@@
Mean Squared Error (MSE) and Peak Signal-to-Noise Ratio (PSNR) (psn, nd) are essential metrics for evaluating image quality and are widely used in AI applications. MSE measures the average squared difference between the pixel values of a reference and processed image, providing a quantifiable representation of distortion. PSNR, derived from MSE, offers a perceptual quality metric expressed in decibels, comparing the maximum signal power with the average squared error. Higher PSNR values indicate high-quality images. While MSE alone may not align with human perception of quality, PSNR provides a standardized measure that accounts for human visual perception.
@@
Wang et al. (Wang and Bovik, 2002) present a universal image quality index, distinct from MSE and PSNR, defined mathematically for independence from specific images, viewing conditions, and observers. This index incorporates correlation coefficient, mean luminance difference, and contrast similarity, offering practical advantages in simplicity and universality. In related work, Wang et al. (Wang et al., 2004) introduce the Structural Similarity (SSIM) index as an alternative to traditional error-sensitive methods, emphasizing structural similarities in luminance, contrast, and structure to align more closely with human visual perception. Furthermore, their Multi-Scale Structural Similarity (MSSIM) approach, as proposed in (Wang et al., 2003), extends SSIM by considering variations in resolution and viewing conditions, providing increased flexibility and demonstrating improved performance.
@@
In a similar vein, Sarnoff’s JND-Metrix (sar, nd) takes into account Just Noticeable Differences (JND) based on the Human Visual System (HVS) principles. Unlike traditional metrics such as PSNR or MSE, the JND-Metrix considers the sensitivity and perception of the HVS to different types of image distortions. By incorporating knowledge of the HVS, including factors like contrast sensitivity, spatial masking, and visual attention, it quantifies the perceptual impact of image distortions more accurately. The method measures the visibility of distortions by estimating JND thresholds for various distortions, indicating the level of distortion perceptually distinguishable from the original image.
@@
Two notable advancements in image quality assessment include Sheikh et al.’s (Sheikh and Bovik, 2006) Visual Information Fidelity (VIF) criterion and Chandler et al.’s (Chandler and Hemami, 2007) Visual Signal-to-Noise Ratio (VSNR) metric. VIF, a full-reference method, evaluates the correlation between image information and visual quality, outperforming traditional metrics in various scenarios. On the other hand, VSNR uniquely considers human visual system properties, including near-threshold and suprathreshold perception, providing a more accurate representation of perceived quality by addressing contrast sensitivity and global precedence.
@@
Lakhani (Lakhani, 2020) and Sabottke et al. (Sabottke and Spieler, 2020) demonstrate that resolution is a critical image quality metric when developing deep learning models for medical imaging and radiology applications. Higher image resolutions can lead to improved AI model performance, especially in detecting specific medical conditions. However, it is essential to strike a balance with computational resources to avoid limitations in the training process. The appropriate image resolution for a given task can vary based on several factors, including the image data type, the specific AI model or algorithm being used, and the application’s requirements. Additionally, blurriness is another factor that can impact the performance of AI models, particularly those tailored for image recognition, object detection, and segmentation tasks. Additionally, blurriness, a significant factor affecting AI model performance, has been extensively studied. Lin et al.’s (Lin et al., 2005) method estimates contrast decrease on edges, while Marziliano et al. (Marziliano et al., 2002) gauge blurriness by analyzing edge spread, providing valuable insights into overall blurriness levels.
@@
Summary: Image quality refers to how accurately an image represents its visual content and cover aspects such as sharpness, clarity, color accuracy, and the absence of artifacts. Commonly used metrics for image quality assessment include MSE, PSNR, SSIM, MSSIM, JND-Metrix, VSNR, and VIF. These metrics provide quantitative measurements of image distortion and quality, allowing practitioners to determine the suitability of image data for AI tasks. Additionally, considering factors like image resolution and blurriness is crucial to balance improved model performance and computational resources.
@@ 4.2.2.2 Speech Quality
Example: Speech quality is important for the optimal functioning of voice-controlled virtual assistants. In a scenario with high speech quality, a user’s command is delivered with clarity, intelligibility, and minimal background noise. This ensures the virtual assistant accurately interprets and executes the request. However, in a low speech quality scenario with distortions and noise, the AI system may struggle to comprehend the command, leading to potential errors in execution.
@@
Metrics in Literature: One commonly used metric is the Mean Opinion Score (MOS), introduced by the ITU-T (International Telecommunication Union – Telecommunication Standardization Sector) (International Telecommunication Union, 2018), involves human listeners rating the perceived quality of speech. MOS helps to understand how well the speech data aligns with human expectations, enabling improvements based on listener feedback. On the other hand, Perceptual Evaluation of Speech Quality (PESQ), introduced by Rix et al. (Rix et al., 2001), quantifies the quality of processed or degraded speech signals objectively. It uses a model that simulates human auditory perception to calculate prediction scores. Segmental Signal-to-Noise Ratio (SNRseg), introduced by Jayant et al. (Jayant and Noll, 1984), is another speech quality metric that gauges the ratio of the clean speech signal to the background noise within short segments. By providing a localized evaluation, SNRseg helps assess the impact of noise on speech intelligibility. It assists in identifying segments of speech that may be affected by noise, leading to targeted noise reduction or enhancement techniques.
@@
Widely used objective metrics for evaluating speech quality and intelligibility are Short-Time Objective Intelligibility (STOI), Perceptual Speech Quality Measure (PSQM), Itakura-Saito Spectral Distortion (IS), and Objective Difference Grade (ODG). STOI, introduced by Taal et al. (Taal et al., 2010), primarily assesses the similarity between clean and degraded speech signals, focusing on how well the degraded speech retains intelligibility compared to the original signal. In contrast, PSQM, introduced by Beerends et al. (Beerends and Stemerdink, 1994), explores various factors affecting speech quality, including distortion, noise, and echo. IS (Itakura and Saito, 1968) quantifies spectral distortion in the frequency domain, helping to understand the impact of different operations on speech quality. On the other hand, ODG, specified in (itu, nd), evaluates the perceived difference in quality between two speech signals, aiding in comparing speech processing algorithms or system configurations. These metrics offer distinct perspectives on speech quality, with STOI and PSQM emphasizing intelligibility and overall quality, IS focusing on spectral distortion, and ODG providing a comparative quality assessment.
@@
Summary: Speech quality refers to the perceived clarity, intelligibility, and fidelity of speech signals. Metrics like Mean Opinion Score (MOS), Perceptual Evaluation of Speech Quality (PESQ), Segmental Signal-to-Noise Ratio (SNRseg), Short-Time Objective Intelligibility (STOI), Perceptual Speech Quality Measure (PSQM), Itakura-Saito Spectral Distortion (IS), and Objective Difference Grade (ODG) are used to assess speech quality objectively and subjectively.
@@ 4.2.2.3 Video Quality
VQM (Video Quality Metric) introduced by Huynh-Thu et al. (Huynh-Thu and Ghanbari, 2008), VMAF (Video Multimethod Assessment Fusion) introduced by Netflix (Blog, 2017), and OPTICOM’s PEVQ (Perceptual Evaluation of Video Quality) (PEV, 2008) are advanced and specialized metrics specifically designed to assess the quality of video data. VQM focuses on replicating human visual perception to evaluate video quality accurately. By analyzing spatial and temporal characteristics of video frames, VQM provides a comprehensive measure of video fidelity, making it invaluable in video compression and transmission applications. VMAF takes a multifaceted approach by combining traditional metrics like PSNR and SSIM with machine learning techniques. VMAF predicts how viewers perceive video quality by leveraging a human-rated dataset, making it highly effective for video streaming, content delivery, and codec research. Meanwhile, PEVQ, standardized by ITU (International Telecommunication Union), offers both objective and subjective evaluations. Its computational model estimates video quality based on visual and temporal features, while human evaluators provide MOS-based subjective assessments. Widely used in video telephony and conferencing systems, PEVQ ensures that video communication meets acceptable quality standards.
@@
Summary: Video quality refers to a video sequence’s overall visual fidelity and perceptual coherence, assessing its accuracy in representing the original content. Objective metrics such as PSNR, SSIM, VIF, and VSNR, which are commonly used for image quality assessment, can be effectively applied to measure video quality by analyzing spatial and temporal characteristics. Additionally, specialized metrics like VQM, VMAF, and PEVQ are specifically designed to address the challenges unique to video data, incorporating human perceptual aspects and machine learning techniques to predict viewer perception.
@@
Preparation of unstructured data for AI applications involves adapting key dimensions typically applied to structured data (4.1). “Correctness” is fundamental, ensuring the accuracy and integrity of content within unstructured data like text, images, audio, and video to maintain AI model reliability. “Feature Relevancy” is crucial for identifying informative elements within unstructured data, aiding pattern recognition and decision-making. “Privacy Leakage” safeguards sensitive information, requiring privacy-preserving techniques. Addressing “Class Imbalance” and “Class Separability” enhances classification tasks in unstructured data by ensuring balanced representation and category distinguishability. Lastly, “Timeliness” is vital, particularly in dynamic data domains, as it ensures AI models stay relevant and up-to-date with evolving data patterns. These dimensions collectively contribute to unstructured data readiness for AI applications.
DRAI Metrics for Unstructured Data

### v2

@@ 3.2.1.5 Bias Indicator
Example: Consider a dataset of job application statements used to train an AI system to evaluate and rank applicants. A bias indicator can be used to analyze the textual data, ensuring fairness and reducing bias. The indicator would examine language and semantic patterns to identify potential biases, such as gender, race, or age-related biases. For instance, if the indicator detects a bias where certain occupations or characteristics are consistently favored or discriminated against, it would flag it for further examination. This enables developers to address and mitigate biases in the dataset before training the AI model. This promotes equal opportunities and minimizes discriminatory outcomes during the evaluation process. The bias indicator plays a crucial role in preparing the textual data, allowing the creation of more objective and unbiased AI systems for job application assessments.
Metrics in Literature: Papakyriakopoulos et al. (2020) propose a robust bias measurement technique using word embeddings and cosine similarity calculations. The method involves defining word pairs to represent different types of discrimination and creating a list of concepts for measuring bias. For variable concepts, which change based on social groups, the bias calculation equation considers the cosine similarity between word pair embeddings and concept embeddings. A modified bias calculation equation is used for non-variable concepts, which remain the same irrespective of social groups. By quantifying the differences in cosine distances, the proposed method comprehensively analyzes bias in different contexts, accounting for variable and non-variable concepts.
Impact on AI: Textual data often contains subtle biases related to gender, race, ethnicity, and other social factors, which can be encoded in language patterns, word associations, and context. If these biases are not identified and addressed before training, an AI model may learn to reinforce these biases, leading to unfair outputs in tasks such as language generation, sentiment analysis, or text classification. This preventive approach leads to ethical and unbiased AI systems.
Summary: A bias indicator is used to quantify and identify potential biases in textual data for AI applications. A bias indicator that uses word embeddings and cosine similarity calculations is one of the most used metrics.
@@ 3.2.2. Multimedia Data
In this subsection, we explore the key aspects of DRAI for image, speech, and video data. We concentrate on quality evaluation metrics for each data type. This exclusive emphasis on quality evaluation metrics is fundamentally crucial in determining the reliability and usability of these types of data for AI.
@@ 3.2.2.1 Image Quality
As a metric to measure the readiness of image data for AI, image quality refers to the degree to which an image accurately represents the visual content it intends to depict. It includes various aspects such as sharpness, color accuracy, clarity, and the absence of artifacts or distortions. Assessing image quality is crucial in AI applications as it directly impacts the reliability and effectiveness of algorithms that rely on visual data. High-quality images provide a solid foundation for AI to extract meaningful features, recognize patterns, and make accurate decisions.
Example: In an AI system designed for autonomous driving, image quality plays a vital role in ensuring the accuracy and reliability of object detection and recognition. High-quality images with clear details and accurate colors allow the AI system to accurately identify pedestrians, vehicles, and traffic signs, enabling it to make precise decisions in real time. Conversely, low-quality images with blurriness or artifacts may lead to misinterpretation of objects or false detection, compromising the safety and efficiency of autonomous driving systems.
@@
Mean Squared Error (MSE) and Peak Signal-to-Noise Ratio (PSNR) (psn, nd) are essential metrics for evaluating image quality and are widely used in AI applications. MSE calculates the average squared difference between the pixel values of a reference and processed image, providing a quantifiable representation of distortion. PSNR, derived from MSE, offers a perceptual quality metric expressed in decibels, comparing the maximum signal power with the average squared error. Higher PSNR values indicate high-quality images. While MSE alone may not align with human perception of quality, PSNR provides a standardized measure that accounts for human visual perception.
@@
Wang and Bovik (2002) present a universal image quality index, distinct from MSE and PSNR, defined mathematically for independence from specific images, viewing conditions, and observers. This index incorporates correlation coefficient, mean luminance difference, and contrast similarity, offering practical advantages in simplicity and universality. In related work, Wang et al. (2004) introduce the Structural Similarity (SSIM) index as an alternative to traditional error-sensitive methods, emphasizing structural similarities in luminance, contrast, and structure to align more closely with human visual perception. Furthermore, their Multi-Scale Structural Similarity (MSSIM) approach, as proposed in (Wang et al., 2003), extends SSIM by considering variations in resolution and viewing conditions, providing increased flexibility and demonstrating improved performance.
@@
Similarly, Sarnoff’s JND-Metrix (sar, nd) takes into account Just Noticeable Differences (JND) based on the Human Visual System (HVS) principles. Unlike traditional metrics such as PSNR or MSE, the JND-Metrix considers the sensitivity and perception of the HVS to different types of image distortions. By incorporating knowledge of the HVS, including factors like contrast sensitivity, spatial masking, and visual attention, the perceptual impact of image distortions is quantified more accurately. JND-Metrix measures the visibility of distortions by estimating JND thresholds for various distortions, indicating the level of distortion perceptually distinguishable from the original image.
@@
Two notable advancements in image quality assessment include Sheikh et al.’s (Sheikh and Bovik, 2006) Visual Information Fidelity (VIF) criterion and Chandler et al.’s (Chandler and Hemami, 2007) Visual Signal-to-Noise Ratio (VSNR) metric. VIF, a full-reference method, evaluates the correlation between image information and visual quality, outperforming traditional metrics in various scenarios. On the other hand, VSNR uniquely considers human visual system properties, including near-threshold and supra-threshold perception. It provides a more accurate representation of perceived quality by addressing contrast sensitivity and global precedence.
The PyTorch Image Quality (PIQ) (Kastryulin et al., 2019) (Kastryulin et al., 2022) provides a diverse array of image quality metrics designed for various assessment requirements. Among full-reference metrics, FSIM (Feature Similarity Index Measure) (Zhang et al., 2011) is significant for its ability to evaluate image quality by analyzing structural and color features. GMSD (Gradient Magnitude Similarity Deviation) (Xue et al., 2014) is also noteworthy which focuses on the gradient magnitudes to assess image quality. In the no-reference category, BRISQUE (Blind/Referenceless Image Spatial Quality Evaluator) (Mittal et al., 2012) stands out as it evaluates quality using natural scene statistics without needing a reference image. This makes it particularly useful in real-world applications. Within the distribution-based category, FID (Frechet Inception Distance) (Heusel et al., 2018) is a key metric that is used for assessing generative models by measuring the similarity between distributions of real and generated images.
@@
Lakhani (2020) and Sabottke and Spieler (2020) demonstrate that resolution is a critical image quality metric when developing deep learning models for medical imaging and radiology applications. Higher image resolutions can lead to improved AI model performance, especially in detecting specific medical conditions. However, it is essential to strike a balance with computational resources to avoid limitations in the training process. The appropriate image resolution for a given task can vary based on several factors, including the image data type, the specific AI model or algorithm used, and the application’s requirements. Additionally, blurriness is another factor that can impact the performance of AI models, particularly those tailored for image recognition, object detection, and segmentation tasks. Additionally, blurriness, a significant factor affecting AI model performance, has been extensively studied. Lin et al.’s (Lin et al., 2005) method estimates contrast decrease on edges, while Marziliano et al. (2002) gauge blurriness by analyzing edge spread, providing valuable insights into overall blurriness levels.
Impact on AI: Measuring image quality impacts AI training by ensuring that the data used to train models is clear, accurate, and free from distortions, which directly influences the model’s ability to learn effectively. High-quality images allow AI models to extract meaningful features and recognize patterns more accurately, leading to better generalization and performance. Conversely, low-quality images with noise, blurriness, or artifacts can introduce errors, reduce the model’s learning efficiency, and lead to inaccurate predictions.
@@
Summary: Image quality refers to how accurately an image represents its visual content and covers aspects such as sharpness, clarity, color accuracy, and the absence of artifacts. Commonly used metrics for image quality assessment include MSE, PSNR, SSIM, MSSIM, JND-Metrix, VSNR, and VIF. These metrics provide quantitative measurements of image distortion and quality. Additionally, considering factors like image resolution and blurriness is crucial to balance improved model performance and computational resources.
@@ 3.2.2.2 Speech Quality
Example: Speech quality is important for the optimal functioning of voice-controlled virtual assistants. In a scenario with high speech quality, a user’s command is delivered with clarity and minimal background noise. This ensures the virtual assistant accurately interprets and executes the request. With low speech-quality data with distortions and noise, an AI system may struggle to comprehend the command, leading to potential errors in execution.
@@
Metrics in Literature: The ITU-T (International Telecommunication Union – Telecommunication Standardization Sector) introduced Mean Opinion Score (MOS) (International Telecommunication Union, 2018), which involves human listeners rating the perceived speech quality. MOS helps to understand how well the speech data aligns with human expectations, enabling improvements based on listener feedback. On the other hand, Perceptual Evaluation of Speech Quality (PESQ), introduced by Rix et al. (2001), objectively quantifies the quality of processed or degraded speech signals. It uses a model that simulates human auditory perception to calculate prediction scores. Segmental Signal-to-Noise Ratio (SNRseg), introduced by Jayant and Noll (1984), is another metric that gauges the ratio of the clean speech signal to the background noise within short segments. By providing a localized evaluation, SNRseg assists in identifying segments of speech that may be affected by noise, leading to targeted noise reduction or enhancement techniques.
@@
Widely used objective metrics for evaluating speech quality and intelligibility are Short-Time Objective Intelligibility (STOI), Perceptual Speech Quality Measure (PSQM), Itakura-Saito Spectral Distortion (IS), and Objective Difference Grade (ODG). STOI, introduced by Taal et al. (2010), primarily assesses the similarity between clean and degraded speech signals, focusing on how well the degraded speech retains intelligibility compared to the original signal. In contrast, PSQM, introduced by Beerends and Stemerdink (1994), explores various factors affecting speech quality, including distortion, noise, and echo. IS (Itakura and Saito, 1968) quantifies spectral distortion in the frequency domain, helping to understand the impact of different operations on speech quality. On the other hand, ODG, specified in (itu, nd), evaluates the perceived difference in quality between two speech signals, aiding in comparing speech processing algorithms or system configurations. These metrics offer distinct perspectives on speech quality, with STOI and PSQM emphasizing intelligibility and overall quality, IS focusing on spectral distortion, and ODG providing a comparative quality assessment.
Impact on AI: Evaluating speech quality can significantly impact AI systems, particularly in speech recognition, synthesis, and processing. By assessing audio quality, AI models can improve their accuracy in recognizing speech across various conditions, including noisy environments and diverse accents. For speech synthesis, quality evaluation leads to natural-sounding and intelligible artificial voices. In audio processing, it enables the development of better noise cancellation and audio enhancement algorithms. Moreover, speech quality assessment improves AI training by providing more accurate data labeling and helping to filter out low-quality samples.
@@
Summary: Speech quality refers to the perceived clarity, intelligibility, and fidelity of speech signals. Metrics such as Mean Opinion Score (MOS), Segmental Signal-to-Noise Ratio (SNRseg), Short-Time Objective Intelligibility (STOI), Perceptual Evaluation of Speech Quality (PESQ), Perceptual Speech Quality Measure (PSQM), Itakura-Saito Spectral Distortion (IS), and Objective Difference Grade (ODG) are used to assess speech quality objectively and subjectively.
@@ 3.2.2.3 Video Quality
VQM (Video Quality Metric) introduced by Huynh-Thu and Ghanbari (2008), VMAF (Video Multimethod Assessment Fusion) introduced by Netflix (Blog, 2017), and OPTICOM’s PEVQ (Perceptual Evaluation of Video Quality) (PEV, 2008) are advanced and specialized metrics specifically designed to assess the quality of video data. VQM focuses on replicating human visual perception to evaluate video quality accurately. By analyzing spatial and temporal characteristics of video frames, VQM provides a comprehensive measure of video fidelity, making it invaluable in video compression and transmission applications. VMAF takes a multifaceted approach by combining traditional metrics like PSNR and SSIM with machine learning techniques. VMAF predicts how viewers perceive video quality by leveraging a human-rated dataset, making it highly effective for video streaming, content delivery, and codec research. Meanwhile, PEVQ, standardized by ITU (International Telecommunication Union), offers both objective and subjective evaluations. Its computational model estimates video quality based on visual and temporal features, while human evaluators provide MOS-based subjective assessments. Widely used in video telephony and conferencing systems, PEVQ ensures that video communication meets acceptable quality standards.
Impact on AI: Evaluating video quality can significantly impact AI training, particularly for models focused on computer vision, video processing, and generation tasks. By assessing video quality, researchers can curate better training datasets, develop more effective data augmentation techniques, create more accurate labels, and improve the assessment of AI-generated content. This process enables the filtering out of low-quality or corrupted videos that could negatively impact model performance while also allowing for the development of better video generation and enhancement algorithms. Furthermore, incorporating human perception metrics into video quality evaluation helps train AI models to produce results that are visually appealing to human viewers.
@@
Summary: Video quality refers to overall visual fidelity and perceptual coherence, assessing its accuracy in representing the original content. Objective metrics such as PSNR, SSIM, VIF, and VSNR, commonly used for image quality assessment, can be effectively applied to measure video quality by analyzing spatial and temporal characteristics. Additionally, specialized metrics like VQM, VMAF, and PEVQ are specifically designed to address the challenges unique to video data, incorporating human perceptual aspects and machine learning techniques to predict viewer perception.
3.2.2.4 Other Metrics for Unstructured Data
@@
Many metrics mentioned in the structured data section (§3.1) are also applicable to measuring readiness of unstructured data. For instance, preparing unstructured data for AI applications involves adapting key dimensions typically applied to structured data (3.1). “Correctness” is fundamental, ensuring the accuracy and integrity of content within unstructured data like text, images, audio, and video to maintain AI model reliability. “Feature Relevancy” is crucial for identifying informative elements within unstructured data, aiding pattern recognition and decision-making. “Privacy Leakage” safeguards sensitive information, requiring privacy-preserving techniques. Addressing “Class Imbalance” and “Class Separability” enhances classification tasks in unstructured data by ensuring balanced representation and category distinguishability. Lastly, “Timeliness” is vital, particularly in dynamic data domains, as it ensures AI models stay relevant and up-to-date with evolving data patterns. Compliance of unstructured data with FAIR principles is a major requirement to make certain the data is ready for AI applications. All these metrics collectively contribute to unstructured data readiness for AI usage.
DRAI Metrics for Unstructured Data In examining unstructured DRAI, this study addresses both textual and multimedia domains. Textual data metrics, including lexical diversity, term importance, readability score, topic coherence, and bias indicators, are discussed with a focus on metrics highlighted in the literature. Simultaneously, multimedia DRAI metrics, such as image, speech, and video quality, are explored in the same context, featuring metrics highlighted in relevant literature.
4. Existing Frameworks
Although not specifically targeting data readiness for AI, numerous frameworks have been developed to evaluate various aspects of data quality. Targeting comprehensive AI readiness evaluation, we have recently developed AI Data Readiness Inspector (AIDRIN) (Hiniduma et al., 2024). We describe here various data quality evaluation frameworks along with AIDRIN.
General data quality toolkits include frameworks like Informatica’s data quality tool (Informatica, nd), an open-source solution for data profiling, cleansing, and monitoring that assesses metrics such as data completeness, accuracy, and reliability. DQLearn (Shrivastava et al., 2020) is another tool in this category, focusing on systematically addressing data quality issues through detection, correction, and custom rule implementation. Additionally, Deequ (Schelter et al., 2018) stands out as a library that allows unit tests for data, facilitating early data quality checks within data pipelines. Moreover, the PIQ (Kastryulin et al., 2019) (Kastryulin et al., 2022) library offers an extensive list of metrics, including both traditional and modern techniques, for evaluating image quality. This library is a valuable resource for researchers and developers working on AI applications, allowing them to select the most appropriate metric based on their specific requirements. PIQ includes metrics categorized as full-reference, no-reference, and distribution-based metrics.
A small number of toolkits target AI data analysis. Among them, the Data Nutrition Project (DNP) Label (Holland et al., 2018) provides a standardized format for presenting essential dataset information, including metadata, variable descriptions, and correlations, aiding in the preparation of quality training data for AI. The Data Readiness Report (Afzal et al., 2020) focuses on generating detailed documentation to assist in data preprocessing for machine learning, offering insights into data quality and readiness across standardized dimensions. IBM’s Data Quality Toolkit (DQT)(et al., 2021) emphasizes automated explanations of data quality across various dimensions, including completeness, feature relevance, label purity, and data fairness, simplifying data preparation and enhancing productivity for AI practitioners.
Domain-specific toolkits are frameworks designed for specific domains or dimensions of DRAI, such as fairness and FAIR compliance. IBM’s AI fairness 360 toolkit (et al., 2018) is an open-source software aimed at detecting and mitigating biases in machine learning models, providing metrics like representation and statistical rates of sensitive attributes. For FAIR compliance, tools like FAIR Cookbook (Rocca-Serra et al., 2023) and FAIRassist (FAIRassist.org, nd) guide users in implementing and measuring FAIR principles, while ESS-DIVE (et al., 2024) evaluates FAIR compliance in earth sciences data repositories.
Toward a comprehensive evaluation of data readiness for AI, we have recently developed AIDRIN (AI Data Readiness Inspector) (Hiniduma et al., 2024). AIDRIN encompasses both traditional data quality assessments and metrics specifically designed for AI readiness, such as data bias, privacy, feature relevance, correlation, and FAIR compliance. The framework allows users to select their assessment criteria and generate intuitive visualizations and reports to enhance the interpretation and usability of results in AI-related tasks.
Despite the increasing interest in evaluating data readiness, there is still a lack of comprehensive tools covering a broad range of metrics to evaluate the readiness of structured and unstructured data for a given AI task. This is a challenging endeavor, and this survey paper will serve as a reference for understanding the available metrics and developing strategies for incorporating them into tools to comprehensively assess data readiness.
5. Pillars of Data Readiness for AI
Based on the knowledge gathered in this survey, we propose a high level taxonomy of metrics. We categorize the AI-ready data assessment metrics into six pillars. These are data quality, understandability and usability, data structure and organization, impact of data on AI, fairness and bias, and data governance (as shown in Fig. 3). These pillars contain a comprehensive set of aspects in data preparation, ensuring that data is prepared and readied to support AI systems effectively, ethically, and efficiently. Each pillar is supported by specific metrics that provide a structured framework for evaluating data readiness in AI contexts.
Figure 3. A high-level view of data readiness metric categories for AI.
Of these categories of metrics, the first four, i.e., data quality, understandability and usability, structure and organization, and data governance, are agnostic to the AI method. They can be applied generically to a broad set of datasets and be used for a wide range of AI applications. In contrast, the last two categories, i.e., the impact of data and fairness, are more specific to the use case and the AI methodology, making them critical for specialized contexts but requiring tailored approaches.
Data quality ensures that the data used to train AI models is accurate, complete, and consistent. High-quality data minimizes the risk of errors in AI outputs, leading to more reliable and trustworthy models. When data quality is compromised, it can result in inaccurate and unstable models. Thus, maintaining rigorous data quality standards is essential for achieving effective and credible AI outcomes. Structured data quality can be evaluated using metrics described in Section 3, such as completeness, correctness, timeliness, mislabeling, multimedia quality.
Data understanding and usability are important for enabling AI systems to interpret and utilize data effectively. This category of metrics emphasize the importance of clear documentation, comprehensive metadata, reusability, and accessible data interfaces. When data is well-understood and easy to use, it accelerates AI model development. FAIR principle compliance metrics[3.1.15] can serve in evaluating data understanding and usability.
Data structure and data organization are important to integrate data into AI workflows seamlessly and efficiently. Adequate number of samples in data and proper data partitioning into training, testing, and validation sets allow for accurate model evaluation. In addition, the data model, i.e., the schema of the data, and data organization, i.e., how the data is stored, also play a role in the speed of AI training. also plays a role in the performance of AI applications.
Data Governance is essential for managing data in a way that is ethical, secure, and compliant with legal standards. This pillar covers key aspects such as data privacy, security, and the ethical use of data, which are necessary for building trust in AI systems. Without proper governance, AI systems risk violating privacy regulations, facing security breaches, and engaging in unethical practices, which can harm public trust and lead to significant legal and reputation-related consequences. Metrics such as privacy leakage[3.1.13] are essential for understanding the extent of potential privacy risks within the data.
The impact of data on AI covers the importance of data content and its relevance to AI applications. Rich and high-impact data that provides meaningful insights is critical for deriving effective AI outcomes by enabling models to make accurate predictions and identify deep patterns. Feature relevance[3.1.5] and data point impact[3.1.10] serve as quantitative measures to assess the value of data for a given AI application.
Fair and unbiased data is a fundamental aspect of ensuring that AI systems produce equitable and unbiased outcomes. This pillar focuses on the representativeness of the data and the absence of biases that could lead to discriminatory practices. Fairness in AI models is not only an ethical concern but also crucial for maintaining public trust in AI technologies. When data used in AI models is biased or unrepresentative, it can result in skewed outcomes that may continue existing inequality and injustice. This undermines the societal benefits that AI promises to deliver. Metrics such as the discrimination (bias) index[3.1.8], class imbalance[3.1.6], and class separability[3.1.7] are crucial for assessing the fairness of data before its use in AI applications.
The six dimensions outlined above can be quantified using specific metrics discussed in this study to evaluate DRAI. While these metrics provide a foundational assessment, additional metrics are required to fully capture the breadth of each dimension. Aggregating the evaluations across the existing and future metrics will result in a comprehensive DRAI assessment that would lead to highly accurate and impactful AI solutions.